OpenStudy (anonymous):

Solve 4x = 12 x ≈ 0.27 x ≈ 0.56 x ≈ 1.79 x ≈ 19.93

OpenStudy (anonymous):

thats 4x or 4x^2

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

4x = 12 x = 3...is the problem written correctly ?

OpenStudy (nincompoop):

it is 4^x = 12 lol

OpenStudy (anonymous):

i think its x^2 which would make root 3 = 1.72

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

take log hehe

OpenStudy (anonymous):

so it will be x log(4) = log(12) x= log(12)/log(4) so calculate now

OpenStudy (anonymous):

log 12 = 1.07 log 4= 0.60

OpenStudy (anonymous):

so x = 1.783333333333

OpenStudy (anonymous):

so its c o_o